OpenStudy (anonymous):

find the minimum and maximum possible areas of a rectangle measuring 2 km by 5 km help please:)!

OpenStudy (jim766):

if the dimensions are given, width = 2 km length = 5 km there is only one possible area A = L * W A = 5 * 2 A = 10

OpenStudy (jim766):

is there more to the question?